OpenStudy (saad560):

What is the value of p? 3+17p=-32-11p

OpenStudy (saad560):

@563blackghost

563blackghost (563blackghost):

This is your question... \(\huge{3+17p = -32 - 11p}\) ?

OpenStudy (saad560):

the value of p

563blackghost (563blackghost):

First we would need to get the numbers without variables to one side so we would subtract 3 from both sides.... \(\huge{(3-3)+17p=(-32-3)-11p}\) What would that equal?

OpenStudy (saad560):

@563blackghost

563blackghost (563blackghost):

Did you find the simplification of.... \(\huge{(3-3)+17p=(-32-3)-11p}\)

OpenStudy (saad560):

17p to -18p

OpenStudy (saad560):

@563blackghost

OpenStudy (saad560):

@awesomedude2078

563blackghost (563blackghost):

Hmmm it would actually be... \(\huge{17p=-35-11p}\) Now we need to get the variable to one side so we would have to add 11p to both sides.... \(\huge{17p+11p=-35(-11p+11p)}\) What does that simplify too?

OpenStudy (saad560):

28p=-35(22p)

OpenStudy (saad560):

0p

OpenStudy (saad560):

instead of 22p

563blackghost (563blackghost):

Correct :) So we have... \(\huge{28p=-35}\) Now we need to isolate the variable...so we would divide by 28 from both sides... \(\Huge\color{Red}{\frac{28p}{28}=\frac{-35}{28}}\) What would p equal?

OpenStudy (saad560):

-35 @563blackghost

OpenStudy (nurali):

p=-35/28

OpenStudy (nurali):

p=?

OpenStudy (saad560):

p=28

563blackghost (563blackghost):

Correct but simplify the fraction :) We list the factors of both 35 and 28... 35: 1, 5, 7, 35 28: 1, 2, 4, 7, 14, 28 What factor do they share?

OpenStudy (saad560):

7

OpenStudy (saad560):

and 1

563blackghost (563blackghost):

So we divide by 7.... \(\huge{\frac{-35}{28} \div \frac{7}{7}=?}\)

OpenStudy (saad560):

1.25

OpenStudy (saad560):

or 5/4

563blackghost (563blackghost):

Actually it will be `-1.25`....So p equals... \(\huge{p=-\frac{5}{4}}\)